OpenStudy (anonymous):

Find the slope of the line that passes through the points (5, 8) and (9, 4).

OpenStudy (anonymous):

\[Slope = \frac{ y2 - y1 }{ x2 - x1 }\]\[Slope = \frac{ 4 - 8 }{ 9 - 5 }\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

3 1 -1 four-thirds

OpenStudy (anonymous):

\[\frac{ 4 - 8 }{ 9 - 5 } = \frac{ -4 }{ 4 }\] What is -4/4 simplified?

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Right! (: